OpenStudy (anonymous):

Differentiate x^2 surd x

OpenStudy (mimi_x3):

\[x^2 \sqrt{x}\]?

OpenStudy (lgbasallote):

\[\Large x^2 \sqrt x \implies x^2 \times x^{1/2} \implies x^{2 + 1/2} \implies x^{5/2}\] does that help?

OpenStudy (lgbasallote):

@Ayeshaafzal you still here?

OpenStudy (anonymous):

But the answer says 5 root x3 over 2

OpenStudy (lgbasallote):

correct...

OpenStudy (lgbasallote):

i only said the equivalent of \[x \sqrt x\] i never said it was the answer...

OpenStudy (lgbasallote):

you have to take the derivative of \[x^{5/2}\] use power rule

OpenStudy (anonymous):

X^3*

OpenStudy (lgbasallote):

you know how to use power rule right?

OpenStudy (anonymous):

No i forgot

OpenStudy (lgbasallote):

\[\huge \frac{d}{dx} (x^n) \implies nx^{n-1}\] remember now?

OpenStudy (anonymous):

Ohhh i get it thnx

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

I am still not getting the answer

OpenStudy (lgbasallote):

did you get \[\huge\frac{d}{dx} (x^{5/2}) \implies \frac 52 x^{\frac 52 - 1}\]

OpenStudy (anonymous):

Are we not gonna times by x^2 which is the value of n

OpenStudy (lgbasallote):

n is 5/2

OpenStudy (lgbasallote):

the form is x^n remember? n is the exponent...in this case it's 5/2

OpenStudy (lgbasallote):

for example.. derivative of x^3 is 3x^(3-1) or 3x^2 derivative of x^4 is 4x^(4-1) or 4x^3 derivative of x^2 is 2x^(2-1) or 2x^1 or simply 2x getting the idea now?

OpenStudy (anonymous):

Got it 5*x^3/2 /2= 5 rootx^3 over2

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

Thank u i am so sorry about my dumbness

OpenStudy (lgbasallote):

haha nahh calculus is hard so dont feel bad :D

OpenStudy (lgbasallote):

anyway you're welcome :D i have to go sleep now so i wsh you luck in these problems

OpenStudy (anonymous):

Have a good night